OpenStudy (anonymous):

marta's checkbook shows that she wrote a check for $987 for building materials. what was the price of materials if sales if sales tax is 5%

OpenStudy (anonymous):

assume price to be x. den our equation would be x+ 5% of x =987 solve n get the answer

OpenStudy (anonymous):

why 5% of x because tax is calculated on the cost price of a product.. and selling price = cost price +tax

OpenStudy (anonymous):

but i need to solve for price before tax

OpenStudy (anonymous):

yeahh u hv assumed x to be price before tax so when u solve that equation u will get x which is price before x

OpenStudy (anonymous):

first u find 5% of 987 that will give u the total amount of sales tax.then u subtract from it from 987. \[5\div 100\times 987=49.35\] \[987-49.35=937.65\] 49.35 is the total sales tax 937.65 is the price for materials